Calculate ΔG° in kJ at 15.4 °C for the reaction of sulfur dioxide gas with chlorine gas to form sulfur oxydichloride gas and chlorine oxide gas. Be sure to balance the reaction using the lowest whole numbers. Use ΔH° and ΔS° to calculate and assume that they are independent of temperature. Use the thermodynamic data provided in the Extras folder on BlackBoard. Report your answer to one place past the decimal.
Answer #1

The reaction that occurs is:

SO2 + 2 Cl2 = Cl2O + SOCl2

Is calculated:

ΔHº = n * ΔHf products - n * ΔHf reactive = (76.08) + (-342.93) - (-296.83) - (2 * 0) = 29.98 kJ / mol

ΔSº = (287.17) + (723.7) - (248.11) - (2 * 222.96) = 316.84 J / mol * K = 0.317 kJ / mol * K

The Gibbs free energy change is calculated for 288.4 K,

ΔG = ΔH - T * ΔS = 29.98 kJ / mol * (288.4 K * 0.317 kJ / mol * K) = - 61.4 kJ / mol

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.